OpenStudy (anonymous):

(150)/1+31e^-.4t =75

OpenStudy (kirbykirby):

\[ \frac{150}{1+31e^{-0.4t}}=75\] ?

OpenStudy (anonymous):

\[75=\frac{ 150 }{ 1+31*e ^{-0.4t} }\]\[1+31e ^{-0.4t}=\frac{ 150 }{ 75 }=2\]\[1=31e ^{-0.4t}\]\[e ^{-0.4t}=1/31=0.032258\]\[\ln \left( e ^{-0.4t} \right)=\ln \left( 0.032258 \right)=-3.43399\]\[-0.4t=-3.43399\]\[t=8.58\]